Prove that if a series of continuous functions converges uniformly, then the sum function is also continuous. 


1
Expert's answer
2021-08-30T16:39:16-0400

We say "\sum^{∞}_{n=i}f_n" converges uniformly to f if "S_N = \sum^N_{n=i} f_n" converges uniformly to f.

Since each "f_n" is continuous, as sum of continuous function is continuous => "S_N" is continuous if "N \in \mathbb{N}" . Since "S_N" converges uniformly to f, given any ε>0 if "N_0 \in \mathbb{N}" such that if "N \u2265N_0"

"|S_N^* -f(x)| < \frac{ε}{3} \;if \;x \in X"

Take any "x_0 \in X" since "S_{N_0}" is continuous "S_0" if "\u03b4>0" such that if "|x_0-x|< \u03b4 = 1" =>

"|S_{N_0}(x_0) -S_{N_0}(x)| < \frac{ε}{3}"

Now

"|f(x_0) -f(x)| ≤ |f(x_0) -S_{N_0}(x_0)| + |S_{N_0}(x_0) -S_{N_0}(x)| + |S_{N_0}(x) -f(x)| \\

≤ \frac{ε}{3} + \frac{ε}{3} + \frac{ε}{3} \\

≤ ε"

"x \in (x_0- δ, x_0+ δ)"

By continuity of "S_{N_0}"

"|S_{N_0}(x) -S_{N_0}(x_0)| < \frac{ε}{3}"

if "|x_0-x| < \u03b4"

If "x_0 \in x" given any ε>0

If S>0 such that

"|x_0-x| < \u03b4" => "|f(x) -f(x_0)|< \u03b5"

Hence f is continuous.